a. Explain the concept of the multiplier, and explain the role of the marginal
propensity to save (MPS) in determining the size of the multiplier.
b. Explain how the size of the multiplier will change when one brings in the role of the marginal tax rate.
c. Using the concepts in parts a and b above, calculate the slope of the AE curve and the size of the multiplier if MPS = 0.20. Then, calculate the revised slope of the AE curve and the multiplier when you know that the imports and the marginal tax rate will reduce the slope of the AE curve by another 0.25.
1
Expert's answer
2014-10-14T12:38:49-0400
a. The marginal propensity to save (MPS) refers to the increase in saving (non-purchase of current goods and services) that results from an increase in income i.e. The marginal propensity to save might be defined as the proportion of each additional dollar of household income that is used for saving. It is also used as an alternative term for the slope of the saving line.
An important implication of Marginal Propensity to Save is measurement of the multiplier. A multiplier measures the magnified change in aggregate product i.e. the gross domestic product, resulting from a change in an autonomous variable (for example, government expenditure, investment expenditures, etc.).
b. The tax multiplier is the negative marginal propensity to consume times one minus the slope of the aggregate expenditures line.
The key feature of the simple tax multiplier that differentiates it from the simple expenditures multiplier is how taxes affect aggregate expenditures. In particular, taxes do not affect aggregate expenditures directly (as do government purchases or investment expenditures). They affect aggregate expenditures indirectly through disposable income and consumption. This gives rise to two important differences compared to the simple expenditures multiplier.
c. If MPS = 0.20, then the size of the multiplier is m = 1/MPS = 1/0.2 = 5. The slope of the AE curve will be 0.2. The positive slope of the aggregate expenditures line is the sum of the marginal propensity to consume (MPC), marginal propensity to invest (MPI), and marginal propensity for government purchases (MPG), less the marginal propensity to import (MPM).
If the imports and the marginal tax rate will reduce the slope of the AE curve by another 0.25, the revised slope of the AE curve will be 0.2 + 0.25 = 0.45. The multiplier will be 1/0.45 = 2.22.